Find the x and y intercepts and graph the equation plotting the intercepts. Please show all of your work. Submit your graph to the Dropbox.-5x+-8y=-40
Answer by nyc_function(2733)

To find the x-intercept, let y = 0 and solve for x.
-5x + -8(0) = -40
-5x + 0 = -40
-5x = -40
x = -40/-5
x = 8
The graph of this equation crosses the x-axis at the point (8,0).

To find the y-intercept, let x = 0 and solve for y.

-5(0) + -8y = -40
0 -8y = -40

-8y = -40
y = -40/-8

y = 5

The graph of this equation crosses the y-axis at the point (0,5).
The x-intercept is x = 8, which is also written as (8,0).
The y-intercept is y = 5, which is also written as (0,5).
Copy and paste the link below to see your graph.
Understand?
http://mathway.com/answer.aspx?p=grap?p=-5x+-8ySMB01-40?p=False?p=False?p=True?p=False?p=True